ホモトピー論シンポジウム

文部省科学研究費総合Ａ（代表森田茂之）による研究集会を催しますので御案内申し上げます。

日時 : 1997 年 11 月 26 日（水） ∼ 11 月 29 日（土）会場 : 高知大学理学部２号館６階理学部第１会議室（６０１号室）

プログラム

11 月 26 日（水）

14:00 ∼ 14:50 {河野明（京大・理） , 佃修一（京大・理） \atop ゲージ群の分類空間について}

15:20 ∼ 16:10 高田功（大島商船高専） Lowest dimensions for immersions of orientable manifolds up to unoriented cobordism

11 月 27 日（木）

10:00 ∼ 10:50 石黒賢士（福岡大・理） Pairings of p-compact groups and H-structures on the classifying spaces of finite loop spaces

11:20 ∼ 12:10 Martin Guest (Univ. of Rochester), Andrzej Kozlowski (富山国際大） , 山口耕平（電気通信大） Configuration space models for the space of maps from a Riemann surface to complex projective space

13:40 ∼ 14:30 Mark Mahowald （Northwestern University） The Hopkins-Miller spectrum EO_{2} and the homotopy groups of spheres.

15:00 ∼ 15:50 河本裕介（広島大・理） Classification of certain three-connected covers with modular cohomology

16:20 ∼ 17:00 {向井純夫（信州大・理） \atop 球面のホモトピー群の元の Moore 空間への持上げについて}

11 月 28 日（金）

10:00 ∼ 10:50 酒井道宏（九州大・数理）

On The Hopf-James invariant over B

11:20 ∼ 12:10 Dominique Arlettaz(Universite de Lausanne), 三村護 (岡山大・理), 中畑弘次 (Universite de Lausanne), 柳田伸顕 (茨城大・教育) The mod 2 cohomology of the linear groups over the ring of integers

13:40 ∼ 14:30 吉村善一（名古屋工大） Quasi KO_{∗}-types of CW -spectra X with KU_{∗}X ≅ F ree ⊕ Z/2^{m} 15:00 ∼ 15:50 栗林勝彦（岡山理科大・理）

Module derivations と有限ループ空間の随伴作用について

16:20 ∼ 17:10 Yongjin Song （Inha University） Application of braid structure

11 月 29 日（土）

10:00 ∼ 10:40 下村克己（高知大・理）

On the homotopy groups of the mapping cone of β_{4} at the prime 2

11:10 ∼ 12:00 {大川哲介（広島工業大） \atop 入射的スペクトラムとコホモロジー局所化}

世話人

高知大学理学部下村克己高知大学理学部逸見 豊
〒 780 高知市曙町２−５−６〒 780 高知市曙町２−５−６
TEL 0888 (44) 0111 （呼出）  TEL 0888 (44) 8272
FAX 0888 (44) 8358  FAX 0888 (44) 8358
E-mail katsumi@math.kochi–u.ac.jp  E-mail hemmi@math.kochi–u.ac.jp

なお、 28 日（金）は懇親会を計画しておりますので、沢山の方々のご参加をお願い申し上げます。会場は大学内を予定しています。